A vertical electric field of magnitude 2.00 x 10^4 N/C exists above the earth' s surface on a day when a thunder storm is brewing . A car with a rectangular size of 6.00 m by 3.00 mis a traveling along a dry gravel rod way sloping downward at 10°. Determine the electric flux through the bottom of the car.


1
Expert's answer
2020-08-02T15:04:09-0400
"\Phi=EA\cos{\theta}\\\Phi=Eab\cos{10}\\\Phi=(20000)(6)(3)\cos{10}=3.55\cdot10^5\ Vm"